Known issues and problems

[On this page, red, bold text indicates action needed.]

Since the construction of the EPD, there have been problems here and there. Usually, we resolve them real time, then move on with our lives actually thinking we'll remember what we did. This page will hopefully solve those "What did we decide about what to do with those QA plot holes" and "Wasn't there a mapping change or something?" moments. This shouldn't be grounds for discussion (reserved for, but instead for descriptions of problems encountered, as well as the solutions and rationale behind them. If you see something that should be here, please put it here and add a reminder to the EPD Google Calendar. This will make all our lives so much easier!

Description and link Decision/Result EPD Calendar reminder date if action required
Dark current oscillating in multiple channels
Revisit after run19 2019-07-13
"Y joints" on fiber bundles breaking Reinforcement design needed 2019-06-17
 Scattered cold tiles in QA plots Replace QT32B daughter board  
 4-tile hole in <ADC> and <TAC> QA plots Daughter card (EQ3,0x1C, DC2) replaced  
 Channel measuring cosmics out of Rx board, but blue in QA plots and not showing any MIP peaks Bad QT32B connector (EQ1-0x10-C) replaced  
 EPD cable weight between Rx boards and QT boards Regularly check cable tension between Rx and QT boards 2019-06-17
Totally dead tile Switched the connector from EQ1-0x1D-B channel 3 to EQ3-0x1D-D channel 1  
Totally dead tile Fixed a bent ground pin  
...  ...